### v1

@@ Abstract
China has a long and rich history, encompassing a vast cultural heritage that includes diverse multimodal information, such as silk patterns, Dunhuang murals, and their associated historical narratives. Cross-modal retrieval plays a pivotal role in understanding and interpreting Chinese cultural heritage by bridging visual and textual modalities to enable accurate text-to-image and image-to-text retrieval. However, despite the growing interest in multimodal research, there is a lack of specialized datasets dedicated to Chinese cultural heritage, limiting the development and evaluation of cross-modal learning models in this domain. To address this gap, we propose a multimodal dataset named CulTi, which contains 5,726 image-text pairs extracted from two series of professional documents, respectively related to ancient Chinese silk and Dunhuang murals. Compared to existing general-domain multimodal datasets, CulTi presents a challenge for cross-modal retrieval: the difficulty of local alignment between intricate decorative motifs and specialized textual descriptions. To address this challenge, we propose LACLIP, a training-free local alignment strategy built upon a fine-tuned Chinese-CLIP. LACLIP enhances the alignment of global textual descriptions with local visual regions by computing weighted similarity scores during inference. Experimental results on CulTi demonstrate that LACLIP significantly outperforms existing models in cross-modal retrieval, particularly in handling fine-grained semantic associations within Chinese cultural heritage.
@@
While existing multimodal methods [19, 4, 6, 17, 18] perform well in general cross-modal retrieval tasks, they face significant challenges when dealing with cultural artifacts such as silk patterns and traditional murals. These artifacts typically exhibit rich artistic elements: Dunhuang murals frequently involve complex visual narratives with multi-object compositions and silk patterns often contain repetitive, delicate motifs. As shown in Fig. 2, the image is a repeating pattern, but the text only describes one unit of the pattern, and even for that single unit, the edge details are ignored, as the text focuses solely on the area within the green frame. Vision-Language Models (VLMs) are deep learning frameworks designed to jointly understand and align visual and textual data [22]. Existing generic VLMs, such as Chinese-CLIP [19], fail to incorporate such information during pre-training and primarily emphasized global alignment, resulting in suboptimal performance in cross-modal retrieval tasks in Chinese Cultural Heritage Documents. To address this limitation, we propose LACLIP, which introduces a Local Visual-Text Alignment module during inference stage of Chinese-CLIP [19]. By employing weighted alignment between local images and global textual descriptions, LACLIP significantly improves cross-modal retrieval performance, as validated by experimental results.
@@ 2.1 Cross-modal Retrieval Datasets
2.2 Cross-modal Retrieval
@@
In contrast to global alignment methods, local alignment approaches aim to capture fine-grained correspondences between specific regions of an image and relevant parts of the text. Multimodal alignment methods show success in tasks like visual question answering (VQA) [1, 8] and fine-grained image-text retrieval. For example, LexVLA [11] has been introduced as an interpretable Visual-Language Alignment framework that learns a unified lexical representation with local alignment and an overuse penalty, while SEA [20] has been proposed as a token-level alignment method that leverages vision-language pre-trained models like CLIP [17] to precisely align visual tokens with the embedding space of Large Language Models (LLMs). However, they typically require extensive pretraining, domain-specific annotations, or tailored architectures, limiting their practicality in the cultural heritage domain.
@@ 3.4 Data Summarization
In summary, we created the first high-quality multimodal Simplified Chinese dataset, which includes 5,726 pairs of Chinese cultural heritage images and corresponding textual description. There are four different categories of data as shown in Fig. 4, including pattern, original textile, cropped pattern, and mural. The overall distribution of the dataset from both data sources is illustrated in Fig. 5. CulTi includes artifacts from different dynasties and artistic styles, highlighting the diversity of the dataset. Within each volume, data of every category—whether original textiles, patterns, cropped patterns, or murals—is independently partitioned into training, testing, and validation subsets in a 7:1:2 ratio, ensuring a balanced and representative distribution for model development and evaluation.
@@ 6 Conclusion
In this work, we introduced CulTi, the first multimodal dataset dedicated to Chinese cultural heritage. CulTi comprises 5,726 high-quality image-text pairs extracted from authoritative publications on ancient Chinese textiles and Dunhuang murals. CulTi fulfill the gap in multimodal cultural area provides valuable resources for advancing research in digital cultural heritage preservation and semantic analysis. We also proposed the LACLIP model, which incorporates a local alignment module into a fine-tuned Chinese-CLIP framework. By emphasizing weighted alignment between local visual features and global textual descriptions, LACLIP significantly enhances cross-modal retrieval performance. In future work, we plan to expand the dataset and explore additional tasks, further refining local alignment techniques to boost performance in even more challenging multimodal scenarios.

### v2

@@ Abstract
China has a long and rich history, encompassing a vast cultural heritage that includes diverse multimodal information, such as silk patterns, Dunhuang murals, and their associated historical narratives. Cross-modal retrieval plays a pivotal role in understanding and interpreting Chinese cultural heritage by bridging visual and textual modalities to enable accurate text-to-image and image-to-text retrieval. However, despite the growing interest in multimodal research, there is a lack of specialized datasets dedicated to Chinese cultural heritage, limiting the development and evaluation of cross-modal learning models in this domain. To address this gap, we propose a multimodal dataset named CulTi, which contains 5,726 image-text pairs extracted from two series of professional documents, respectively related to ancient Chinese silk and Dunhuang murals. Compared to existing general-domain multimodal datasets, CulTi presents a challenge for cross-modal retrieval: the difficulty of local alignment between intricate decorative motifs and specialized textual descriptions. To address this challenge, we propose LACLIP, a training-free local alignment strategy built upon a fine-tuned Chinese-CLIP. LACLIP enhances the alignment of global textual descriptions with local visual regions by computing weighted similarity scores during inference. Experimental results on CulTi demonstrate that LACLIP significantly outperforms existing models in cross-modal retrieval, particularly in handling fine-grained semantic associations within Chinese cultural heritage. The dataset and source code are available at: https://github.com/yyyjjy/CulTi
@@
While existing multimodal methods [21, 4, 6, 18, 20, 16] perform well in general cross-modal retrieval tasks, they face significant challenges when dealing with cultural artifacts such as silk patterns and traditional murals. These artifacts typically exhibit rich artistic elements: Dunhuang murals frequently involve complex visual narratives with multi-object compositions and silk patterns often contain repetitive, delicate motifs. As shown in Fig. 2, the image is a repeating pattern, but the text only describes one unit of the pattern, and even for that single unit, the edge details are ignored, as the text focuses solely on the area within the green frame. Vision-Language Models (VLMs) are deep learning frameworks designed to jointly understand and align visual and textual data [24]. Existing generic VLMs, such as Chinese-CLIP [21], fail to incorporate such information during pre-training and primarily emphasized global alignment, resulting in suboptimal performance in cross-modal retrieval tasks in Chinese Cultural Heritage Documents. To address this limitation, we propose LACLIP, which introduces a Local Visual-Text Alignment module during inference stage of Chinese-CLIP [21]. By employing weighted alignment between local images and global textual descriptions, LACLIP significantly improves cross-modal retrieval performance, as validated by experimental results.
@@ 2.1 Cross-modal Retrieval Datasets
2.2 Cross-modal Retrieval Models
@@
In contrast to global alignment methods, local alignment approaches aim to capture fine-grained correspondences between specific regions of an image and relevant parts of the text. Multimodal alignment methods show success in tasks like visual question answering (VQA) [1, 8] and fine-grained image-text retrieval [19]. For example, LexVLA [11] has been introduced as an interpretable Visual-Language Alignment framework that learns a unified lexical representation with local alignment and an overuse penalty, while SEA [22] has been proposed as a token-level alignment method that leverages vision-language pre-trained models like CLIP [18] to precisely align visual tokens with the embedding space of Large Language Models (LLMs). However, they typically require extensive pretraining, domain-specific annotations, or tailored architectures, limiting their practicality in the cultural heritage domain.
@@ 3.4 Data Summarization
In summary, we presented CulTi, a high-quality multimodal dataset focused on Chinese cultural heritage, consisting of 5,726 image-text pairs in simplified Chinese. There are four different categories of data as shown in Fig. 4, including pattern, original textile, cropped pattern, and mural. The overall distribution of the dataset from both data sources is illustrated in Fig. 5. CulTi includes artifacts from different dynasties and artistic styles, highlighting the diversity of the dataset. Within each volume, data of every category—whether original textiles, patterns, cropped patterns, or murals—is independently partitioned into training, testing, and validation subsets in a 7:1:2 ratio, ensuring a balanced and representative distribution for model development and evaluation.
@@ 6 Conclusion
In this work, we introduced CulTi, the first multimodal dataset dedicated to Chinese cultural heritage. CulTi comprises 5,726 high-quality image-text pairs extracted from authoritative publications on ancient Chinese textiles and Dunhuang murals. CulTi fulfill the gap in multimodal cultural area provides valuable resources for advancing research in digital cultural heritage preservation and semantic analysis. We also proposed the LACLIP model, which incorporates a local alignment module into a fine-tuned Chinese-CLIP framework. By emphasizing weighted alignment between local visual features and global textual descriptions, LACLIP significantly enhances cross-modal retrieval performance. In future work, we plan to expand the dataset and explore additional tasks, further refining local alignment techniques to boost performance in even more challenging multimodal scenarios. {credits}
6.0.1 Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62276258), the XJTLU Research Development Fund (REF-22-01-002), the Key Technology R&D Program of Ningbo (2023Z143), and the Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004).